Solve by the Method of Undetermined Coefficients. 1. - 3y - 4y = 3e2x (ans. y = C1e4x + cze* - e2x) 2. - 4y = 4e3x (ans.
Answer #1

--) M:4,-1. PL = 1 in ean Goven y- uy = 4034 Guven y - By- 4y = 3e? Auxiliary equation m?-3m-450 =m-4m+m-4 = 0 = (-4) + (Now the particular integral Pl= I 4 en 4.03h 110 02-4 ean Since I ean. f(0) i if f(a) 40. f(a) en = 32 4 3-4 This Pl= 4 0²-4- [o-Cap?+30) + (20++30)*-- ] *? + 3 lenin 20 + 3D +1 Since I = l-n+m²+ (-43)+.. In Now Pl= (n - (199+30) m? + (404+90°+ 100

Now denn = Renn -10 302-1 Thus. 3-) - 3 (30+) Co i [3 D semn + Roon] renn = renn 20+50 +1 -3 to [scosnt eon] Thus Pl= n²-6n +Now the particular integred! P1 = 2n D²+D-2 +1-(00 I-D²+D Since I = 1+ n²tn3+n4+... - this + PI - (0²403 : Jn 04440-] M + 0 +This 5 Given yu+ 2y +9y= 4e Coran. ; 4(0)=1, y(o)=0 Auxiliary equation m? + 2m +5=0 =) mz -2 254-20 2 = -12 thus the compAs we know for Coran = и 뽑 senan D*+4? ht Thus А: 1 \- Хот де де и Casan 6. - и и1 от и Theis y (n) = (FAPI м) 4 **[, а х +thus y(o)=0 then [0+2C,] + [9 +0]+ 1[+0] + +0=0 => 262 +6,=0 => 2/4=(, = 1 - C2 = 2 This + g(n)= em cas2n the lonan the seni